# ROTI-GCV: Generalized Cross-Validation for right-ROTationally Invariant Data

### v1

@@ A.8 Proof of Theorem 3
Most computations are analogous to the proof of Theorem 2. We first compute $\mathbb{E}[\widetilde{\mathbf{X}}^{\top}\widetilde{\mathbf{X}}\mid\mathbf{X},\mathbf{y}]$ using [33, Lemma 4]. We find that it is equal to $\sum_{i\in\mathcal{J}_{c}}(\mathfrak{d}_{i}^{2}-\mathfrak{d}_{\mathsf{bulk}}^{2})\mathbf{o}_{i}\mathbf{o}_{i}^{\top}+\mathfrak{d}_{\mathsf{bulk}}^{2}\mathbf{I}$. Call this quantity $\mathbf{C}$ for convenience. We then move to computing the risk.
@@
On the other hand, the second term is order $1$ and hence vanishes when divided by $n$. The convergence is then uniform because the function is monotonic in $\lambda$. It is almost sure through, for instance, using [6, Lemma C.3]. Note that the form of $\mathcal{V}_{\mathbf{X}}$ contains an extra
$\frac{\sigma^{2}}{n}\sum_{i\in\mathcal{J}_{c}}\frac{D_{ii}^{2}}{D_{ii^{2}}+\lambda}\mathfrak{d}_{i}^{2}$
@@
    The first term is treated equivalently to the bias in the ridge case in Theorem 2, where, as in the variance computation above, the $\mathfrak{d}_{\mathsf{bulk}}^{2}$ can be handled. This produces uniform convergence to its expectation. The second term is $O(|\mathfrak{J}_{c}|)$ and hence vanishes when divided by $n$. It does so uniformly because it is monotone increasing in $\lambda$.
@@
Recall we are dividing everything by $n$. Hence if we show that $|\mathbf{q}_{i}^{\top}\boldsymbol{\epsilon}|/\sqrt{n}\xrightarrow{a.s.}0$, then we will be done with a union bound. We rewrite $|\mathbf{q}_{i}^{\top}\boldsymbol{\epsilon}|=\boldsymbol{\epsilon}^{\top}(\mathbf{q}_{i}\mathbf{q}_{i}^{\top})\boldsymbol{\epsilon}$ and can now apply [6, Lemma C.3] to find
@@ A.9 Proof of Lemma 5
The following is taken from [27]. We note that in this work, the errors are assumed Gaussian. We will mark which ones require this and how we believe they can be relaxed. First, the consistency of $\alpha_{i}$ is immediate from the first result of [27, (40)], where we note that the proof does not require Gaussianity of the errors. The one step in the proof where care must be made is in controlling
@@
but we note that this can be done through [6, Lemma C.3].
Then the reduced model is right-rotationally invariant by [27, p. 58, (124)]. Showing that the reduced model is right-rotationally invariant requires Gaussianity and is what leads to the added assumption of Lemma 5. Without the Gaussian assumption, the errors in the reduced model will not be i.i.d. However, in general we only require that certain inner products concentrate. We believe this will still hold in the reduced model when one explicitly tracks the form of the error.
@@ A.10 Proof of Corollary 6
Note that $T_{1}=\mathcal{V}_{\mathbf{X}}$. First, note that $\hat{A}$ and $\hat{\sigma}^{2}/n$ are both almost surely bounded for sufficiently large $n$. This follows from the fact that the estimators are strongly consistent for bounded quantities. Note that the fact that we allow $o(n)$ fluctuations on $\hat{\mathfrak{d}}_{i}$ is allowed because there are finitely many of these, and we later divide by $n$. We can thus uniformly control $T_{1}-T_{3}$ and then $T_{3}-T_{2}$, producing uniform convergence.
@@ A.11 Weak Convergence of Spectral Measure
    Almost surely, the diagonal entries of $\mathbf{D}$ converge in the Wasserstein-2 ($W_{2}$) distance to a measure $\mathsf{D}$, where $\mathsf{D}^{2}$ has non-zero mean and compact support $\mathrm{supp}(\mathsf{D}^{2})\subseteq[0,\infty)$. See Appendix B.3 for details on Wasserstein-2 convergence.
@@
This is then similar in flavor to [31, Theorem 4.1].
@@ B.1 Experimental details
B.2 LOOCV Intractability
@@
B.3 Wasserstein-2 Convergence
The following section is taken from [27, Definition 2.2]. For a matrix $(\mathbf{v}_{1},\dots,\mathbf{v}_{k})=(v_{i,1},\dots,v_{i,k})_{i=1}^{n}\in\mathbb{R}^{n\times k}$ and a random vector $(\mathsf{V}_{1},\dots,\mathsf{V}_{k})$, we write
$(\mathbf{v}_{1},\dots,\mathbf{v}_{k})\xrightarrow{W_{2}}(\mathsf{V}1_{,}\dots,\mathsf{V}_{k})$
to mean that the empirical distribution of the columns of $\mathbf{v}_{1},\dots,\mathbf{v}_{k})$ converge to $\mathsf{V}_{1},\dots,\mathsf{V}_{k}$ in Wasserstein-2 distance. This means that for any continuous function $f:\mathbb{R}^{k}\to\mathbb{R}$ satisfying
$|f(x,\dots,x_{k})|\leq C\left(1+\|(x_{1},\dots,x_{k})\|^{2}\right)$ (21)
for some $C>0$, we have
$\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}f(v_{i,1},\dots,v_{i,k})=\mathbb{E}\left[f(\mathsf{V}_{1},\dots,\mathsf{V}_{k})\right]$
where $\mathbb{E}\left[\|(\mathsf{V}_{1},\dots,\mathsf{V}_{k})\|^{2}\right]<\infty$.

### v2

@@ A.8 Proof of Theorem 3
Most computations are analogous to the proof of Theorem 2. We first compute $\mathbb{E}[\widetilde{\mathbf{X}}^{\top}\widetilde{\mathbf{X}}\mid\mathbf{X},\mathbf{y}]$ using [31, Lemma 4]. We find that it is equal to $\sum_{i\in\mathcal{J}_{c}}(\mathfrak{d}_{i}^{2}-\mathfrak{d}_{\mathsf{bulk}}^{2})\mathbf{o}_{i}\mathbf{o}_{i}^{\top}+\mathfrak{d}_{\mathsf{bulk}}^{2}\mathbf{I}$. Call this quantity $\mathbf{C}$ for convenience. We then move to computing the risk.
@@
On the other hand, the second term is order $1$ and hence vanishes when divided by $n$. The convergence is then uniform because the function is monotonic in $\lambda$. It is almost sure through, for instance, using [7, Lemma C.3]. Note that the form of $\mathcal{V}_{\mathbf{X}}$ contains an extra
$\frac{\sigma^{2}}{n}\sum_{i\in\mathcal{J}_{c}}\frac{D_{ii}^{2}}{D^{2}_{ii}+\lambda}\mathfrak{d}_{i}^{2}$
@@
    The first term is treated equivalently to the bias in the ridge case in Theorem 2, where, as in the variance computation above, the $\mathfrak{d}_{\mathsf{bulk}}^{2}$ can be handled. This produces uniform convergence to its expectation. The second term is $O(|\mathcal{J}_{c}|)$ and hence vanishes when divided by $n$. It does so uniformly because it is monotone increasing in $\lambda$.
@@
Recall we are dividing everything by $n$. Hence if we show that $|\mathbf{q}_{i}^{\top}\boldsymbol{\epsilon}|/\sqrt{n}\xrightarrow{a.s.}0$, then we will be done with a union bound. We rewrite $|\mathbf{q}_{i}^{\top}\boldsymbol{\epsilon}|=\boldsymbol{\epsilon}^{\top}(\mathbf{q}_{i}\mathbf{q}_{i}^{\top})\boldsymbol{\epsilon}$ and can now apply [7, Lemma C.3] to find
@@ A.9 Proof of Lemma 4
The following is taken from [23]. We note that in this work, the errors are assumed Gaussian. We will mark which ones require this and how we believe they can be relaxed. First, the consistency of $\alpha_{i}$ is immediate from the first result of [23, (40)], where we note that the proof does not require Gaussianity of the errors. The one step in the proof where care must be made is in controlling
@@
but we note that this can be done through [7, Lemma C.3].
Then the reduced model is right-rotationally invariant by [23, p. 58, (124)]. Showing that the reduced model is right-rotationally invariant requires Gaussianity and is what leads to the added assumption of Lemma 4. Without the Gaussian assumption, the errors in the reduced model will not be i.i.d. However, in general we only require that certain inner products concentrate. We believe this will still hold in the reduced model when one explicitly tracks the form of the error.
@@ A.10 Proof of Corollary 5
Note that $T_{1}=\mathcal{V}_{\mathbf{X}}$. First, note that $\hat{A}$ and $\hat{\sigma}^{2}/n$ are both almost surely bounded for sufficiently large $n$. This follows from the fact that the estimators are strongly consistent for bounded quantities. We can thus uniformly control $T_{1}-T_{3}$ and then $T_{3}-T_{2}$, producing uniform convergence.
@@ A.11 Weak Convergence of Spectral Measure
    Almost surely, the diagonal entries of $\mathbf{D}$ converge in the Wasserstein-2 ($W_{2}$) distance to a measure $\mathsf{D}$, where $\mathsf{D}^{2}$ has non-zero mean and compact support $\mathrm{supp}(\mathsf{D}^{2})\subseteq[0,\infty)$. See Appendix LABEL:app:w2 for details on Wasserstein-2 convergence.
@@
This is then similar in flavor to [29, Theorem 4.1].
@@ B.1 Experimental details
B.2 Semi-real experimental details
Figure 3: TR (Tuned Risk): test risk attained by the value of $\lambda$ chosen through the given CV method. ER (Estimated Risk): test risk estimated by the CV method. MR (MinRisk): attainable value of out-of-sample risk. Plot format same as in main manuscript. Each plot shows 10 resamplings of noise. Dashed blue line shows average actual MSE (over resamplings). Black line shows average CV risk curve for that CV method. Standard deviation of all quantities shown in parentheses.
  1. 1.
    There is one row in $X$ for each minute. Each row is the residualized return over the last 30 minutes, so consecutive rows are heavily correlated, since they share 29 minutes of returns. $X$ has 493 rows and 493 columns ($n=p=493$).
  2. 2.
    We generate $\beta$ uniformly on the sphere of radius $r^{2}\sqrt{n}$, and the training labels $y=X\beta+\epsilon$, where $\epsilon_{i}\sim N(0,\sigma^{2})$.
  3. 3.
    We plot the predicted loss curve given by each cross-validation metric
  4. 4.
    We compute the actual loss curve over the test data.
  5. 5.
    While in the other plots in the main manuscript, we only resampled $\epsilon$, the choice of $\beta$ affects the loss curves more here. We resample $\beta$ four times (referring to Trials 1 through 4 in Figure 3), where we resample $\epsilon$ ten times per trial. In the main manuscript, only one trial was shown.
B.3 Proof of right-rotational invariance
Both the equicorrelated and autocorrelated cases can be written as $\boldsymbol{\Sigma}^{1/2}\mathbf{G}$, where $\mathbf{G}$ has i.i.d. standard Gaussian entries; the rotational invariance of the Gaussian then implies the right-rotational variance of $\boldsymbol{\Sigma}^{1/2}\mathbf{G}$ in these cases.
$t$-distributed data: A multivariate $t$ distribution with $\nu$ degrees of freedom can be generated by first sampling $\mathbf{y}\sim\mathsf{N}(\mathbf{0},\mathbf{I}_{n})$ and $u\sim\chi_{\nu}$; then $\mathbf{x}=\mathbf{y}/\sqrt{u/\nu}$; the rotational invariance of the Gaussian implies that this distribution is rotationally invariant. This proves rotational invariance for a single sample; stacking multiple independent samples into a matrix retains the rotational invariance.
Products of Gaussian matrices: follows from rotational invariance of the Gaussian.
Matrix normals: Recall that $\boldsymbol{\Sigma}^{\mathrm{col}}$ is drawn from an inverse Wishart distribution with identity scale and $(1+\delta)p$ degrees of freedom; this can be generated by $(\mathbf{G}_{2}^{\top}\mathbf{G}_{2})^{-1}$, where $\mathbf{G}_{2}\in\mathbb{R}^{(1+\delta)p\times p}$ has i.i.d. standard Gaussian entries (since the inverse Wishart has identity scale) [11, Appendix A]. A matrix normal with row covariance $\boldsymbol{\Sigma}^{\text{row}}$ and column covariance $\boldsymbol{\Sigma}^{\text{col}}=(\mathbf{G}_{2}^{\top}\mathbf{G}_{2})^{-1}$ can then be generated by $(\boldsymbol{\Sigma}^{\text{row}})^{1/2}\mathbf{G}(\mathbf{G}_{2}^{\top}\mathbf{G}_{2})^{-1/2}$ (with $\mathbf{G}$ again having i.i.d. Gaussian entries and being independent of $\mathbf{G}_{2}$). The claim then follows from the following. For any fixed $\mathbf{O}\in\mathbb{O}_{p}$,
$(\boldsymbol{\Sigma}^{\text{row}})^{1/2}\mathbf{G}(\mathbf{G}_{2}^{\top}\mathbf{G}_{2})^{-1/2}\mathbf{O}$ =$\stackrel{{\scriptstyle d}}{{=}}(\boldsymbol{\Sigma}^{\text{row}})^{1/2}\mathbf{G}\mathbf{O}(\mathbf{O}\mathbf{G}_{2}^{\top}\mathbf{G}_{2}\mathbf{O}^{\top})^{-1/2}\mathbf{O}$
$=(\boldsymbol{\Sigma}^{\text{row}})^{1/2}\mathbf{G}(\mathbf{G}_{2}^{\top}\mathbf{G}_{2})^{-1/2}$
as required.
Spiked matrices: It suffices to show that for any fixed $\mathbf{O}$ that $(\lambda\mathbf{V}\mathbf{W}^{\top},\mathbf{G})\stackrel{{\scriptstyle d}}{{=}}(\lambda\mathbf{V}\mathbf{W}^{\top}\mathbf{O},\mathbf{G}\mathbf{O})$. Since the first consists of an independent pair, it suffices to show that $\mathbf{V}\mathbf{W}^{\top}\stackrel{{\scriptstyle d}}{{=}}\mathbf{V}\mathbf{W}^{\top}\mathbf{O}$, $\mathbf{G}\stackrel{{\scriptstyle d}}{{=}}\mathbf{G}\mathbf{O}$ and that $\lambda\mathbf{V}\mathbf{W}^{\top}\mathbf{O}$ is independent of $\mathbf{G}\mathbf{O}$. The last two statements are automatic from the rotational invariance of the Gaussian and the independence of the $\mathbf{V}\mathbf{W}^{\top}$ and $\mathbf{G}$. The first can be seen from the fact that $\mathbf{W}=\mathbf{W^{\prime}}\mathbf{P}$, where $\mathbf{P}\in\mathbb{R}^{p\times r}$ has $\mathbf{P}_{ij}=\mathds{1}(i=j)$ and $\mathbf{W^{\prime}}\in\mathbb{R}^{p\times p}$ is Haar. Then $\mathbf{V}\mathbf{W}^{\top}\mathbf{O}=\mathbf{V}\mathbf{P}^{\top}(\mathbf{W^{\prime}})^{\top}\mathbf{O}\stackrel{{\scriptstyle d}}{{=}}\mathbf{V}\mathbf{P}^{\top}(\mathbf{W^{\prime}})^{\top}=\mathbf{V}\mathbf{W}^{\top}$ where the distributional equality follows from the rotational invariance of Haar matrices.
B.4 Uniform convergence
Suppose one has a set of (possibly random) functions $f_{n}$ satisfy, for some deterministic $f$,
$\sup_{x\in[x_{1},x_{2}]}|f_{n}(x)-f(x)|\xrightarrow{a.s.}0.$ (29)
Define
$x_{\mathsf{cv},n}=\operatorname*{arg\;min}_{x\in[x_{1},x_{2}]}f_{n}(x)\qquad x_{*}=\operatorname*{arg\;min}_{x\in[x_{1},x_{2}]}f(x).$
Then
$\left|f\left(x_{\mathsf{cv},n}\right)-f(x_{*})\right|\to 0.$ (30)
Here the random functions $f$ represent cross validation metrics which uniformly converge to the true asymptotic out-of-sample risk $f$. This result then shows that the minimizer chosen by the cross-validation metric tends towards the asymptotically optimal value. Note that this result cannot be obtained with only pointwise convergence.
Proof.
Note $f(x_{\mathsf{cv},n})-f(x_{*})\geq 0$ always. Hence it suffices to show an upper bound which tends to zero.
$f(x_{\mathsf{cv},n})-f(x_{*})$ $=f(x_{\mathsf{cv},n})-f_{n}(x_{\mathsf{cv},n})+\underbrace{f_{n}(x_{\mathsf{cv},n})-f_{n}(x_{*})}_{\leq 0}+f_{n}(x_{*})-f(x_{*})$
$\leq|f(x_{\mathsf{cv},n})-f_{n}(x_{\mathsf{cv},n})|+|f_{n}(x_{*})-f(x_{*})|$
$\leq 2\sup_{x\in[x_{1},x_{2}]}|f_{n}(x)-f(x)|\to 0.$
as required. ∎
B.5 Distribution notation
  • $\mathsf{InvWishart}(\mathbf{\Psi},\nu)$ denotes an inverse Wishart distribution [11, Appendix A] with scale matrix $\boldsymbol{\Psi}$ and degrees of freedom $\nu$.
  • $\mathsf{MN}(\boldsymbol{\mu},\boldsymbol{\Sigma}^{\text{row}},\boldsymbol{\Sigma}^{\text{col}})$ denotes a matrix normal with among-row covariance $\boldsymbol{\Sigma}^{\text{row}}$ and among-column covariance $\boldsymbol{\Sigma}^{\text{col}}$.
B.6 LOOCV Intractability
@@
B.7 Algorithm Complexity
Assuming matrix multiplications and inversions cost $O(p^{3})$, ROTI-GCV takes $O(p^{3})$ time. The entire process takes only a constant factor more time than computing a single ridge solution, as the fitting process implicitly requires fitting multiple ridge solutions.